Write an equation of the line passing through (–8, –3) and perpendicular to y =1/4x+2 . Write your answer in standard form Ax + By = C, A > 0
Answer by lwsshak3(11628)

Write an equation of the line passing through (–8, –3) and perpendicular to y =1/4x+2 . Write your answer in standard form Ax + By = C, A > 0
***
equation of a line: y=mx+b, m=slope. b=y-intercept
y =1/4x+2
slope=1/4
slope of line perpendicular to this line=-4 (negative reciprocal)
equation:
y=-4x+b
solve for b using given coordinates (-8,-3) on this line.
-3=-4*-8+b
b=-35
equation: y=-4x-35
or in standard form:
4x+y=-35